Find two consecutive positive odd integers whose sum is 76.
Answer by ikleyn(52776)

.
x + (x+2) = 76,

2x + 2 = 76,

2x = 76 -2,

2x = 74,

x = 37.

The two consecutive odd integers are 37 and 39.